# Twisted Superfields

### v8

@@ 2 Dynamical Symmetry Breaking
$T_{23}^{Y^{\prime}}=T_{22}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(-1,-1,-1,1,1,0\right)$ (2.15)
$T_{21}^{Y^{\prime}}=T_{1F}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(2,2,2,-1,-1,-2\right)$
Either version of $T_{mF}^{Y}$ can be used when acting on gaugino or adjoint fields since $T_{+}^{0}$ commutes with those. The $A_{\mu}^{Y^{\prime}}$ group structures when acting on gauginos and adjoint fermions is the same as $T_{23}^{Y^{\prime}}$.
It will be seen below that the $T_{mF}^{Y}$ matrices have the correct form for their gauge boson $A_{\mu}^{Y}$ to be identified as the U(1) weak hypercharge field of the Standard Model with $-{\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}g_{Y}$ identified as the weak hypercharge coupling. The $A_{\mu}^{Y^{\prime}}$ gauge field with its coupling $g_{Y^{\prime}}$ is a second U(1) gauge boson in this model that remains massless at the unification scale. Due to eq (2.10), there is no mixing between the $A_{\mu}^{Y}$ and $A_{\mu}^{Y^{\prime}}$ gauge fields. Below the unification scale, the couplings for the SU(3), SU(2), $Y$ and $Y^{\prime}$ groups run differently, so they are denoted by $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$.
@@
The angle $\theta_{Z}$ is chosen to make $(T_{21}^{\gamma})_{44}=(T_{22}^{\gamma})_{55}=0$, so that the photon gets no mass from $\bar{\phi}_{21}$, $\bar{\phi}_{22}$ or $\bar{\varphi}_{2}$.
The angle $\phi_{Z}$ is chosen to make $(T_{21}^{Z^{\prime}})_{44}=0$. The resulting angles are given by:
@@
$\tan\phi_{Z}$ $=\tfrac{2}{\sqrt{10}}\cos\theta_{Z}g_{Y^{\prime}}/g_{2}.$
In section 4, it will be argued that $g_{Z^{\prime}}$ will be driven to a very small value from an effective anomaly. As a result, the angle $\phi_{Z}$ is very small, and $\theta_{Z}$ is very close to the Weinberg angle $\theta_{W}$ of the Standard Model.
@@
In the above labelling, lower numerical indices are fundamental-representation indices for the unbroken SU(3) group (the strong interaction). Fermions with a “W” index interact with the $W$ boson (as members of an isodoublet). Based on the magnitude of their electric charges, up-type quarks, down-type quarks, charged leptons, and neutrinos are labelled with $u,d,e,\nu$. All of the fermion fields are 2-component Weyl fermions with a lower, undotted spin index. In the convention of [6] (which is also the Wess/Bagger & Bilal convention of [14]), a Weyl fermion with a lower undotted index corresponds to a right-chiral fermion that vanishes when acted on by $1-\gamma_{5}$ (see appendix A of [15]).
In that convention, the $u$ and $d$ fields are right-chiral fermions with electric charges of ${\tfrac{2}{3}}$ and $-{\tfrac{1}{3}}$, respectively, so they are mapped to right-chiral quarks. The fields $\tilde{u}$ and $\tilde{d}$ are right-chiral fermions with electric charges of $-{\tfrac{2}{3}}$ and ${\tfrac{1}{3}}$, respectively, so they are mapped to Hermitian conjugates of left-chiral quarks. One way that this model differs from the Standard Model is that some of the right-chiral quarks have a “$W$” index so they interact with the $W$ boson, while some of the left-chiral quarks lack that index so they do not interact with the $W$ boson. That difference is discussed in the next section.
@@ 3.1 Confinement
$-\sqrt{2}g_{3}i\int d^{4}x\rm{tr}\left(\tilde{\chi}\left\langle\varphi_{1}^{\dagger}\right\rangle\lambda-\tilde{\lambda}\left\langle\varphi_{1}^{\dagger}\right\rangle\chi\right)+h.c.,$ (3.2)
@@
There are a couple of additional quantum effects pertinent to this section. For supersymmetric theories, instantons can generate nonperturbative low-energy effective superpotential terms. The Appendix proposes that a similar effect occurs for this theory. To accommodate these terms, the following vevs that are zero classically acquire small but nonzero vevs quantum mechanically:
$\bar{\tilde{\phi}}_{21},\bar{\tilde{\phi}}_{22}\neq 0\textrm{ quantum mechanically}.$ (3.3)
A quantum superpotential term proportional to $1/\bar{\tilde{\phi}}_{21}$ (see eq (A.13)) sets the scale for masses of the heavy right-handed neutrinos discussed later in this section.
As discussed in the Appendix, the quantum-generated superpotential terms also try to reduce the value of $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$. To accomplish this in the presence of eq (3.1), the quantum vacuum modifies $\left\langle\varphi_{2}\right\rangle$ of eq (2.6) to the following:
$\left\langle\varphi_{2}\right\rangle=\frac{\bar{\varphi}_{2}}{\sqrt{2\left(1+c^{2}\right)}}\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{0}&{1}\\ {0}&{c}&{0}\end{array}\right),$ (3.4)
where the quantum-generated $c$ is much smaller than 1. The square of the above vev has a nonvanishing trace that can reduce $\left\langle{\rm tr}\left(\varphi_{1}^{2}\right)\right\rangle+\left\langle{\rm tr}\left(\varphi_{2}^{2}\right)\right\rangle$.
@@ 3.2 Observed boson masses
Phenomenologically, the mass of the $W$ boson determines the vev $\bar{\phi}_{21}^{2}$, with the running coupling $g_{2}^{2}$ evaluated at the $W$ boson mass scale.
The $Z$ boson mass in this model is also primarily determined by $\bar{\phi}_{21}^{2}$. Due to the form of eq (2.16), the $Z$ boson mass in this model differs slightly from the Standard Model expression. The mass of the $Z$ boson is:
@@
In this model, the scalar vev $\bar{\phi}_{21}$ also generates most of the mass of the observed Brout-Englert-Higgs boson, through the d-term part of the scalar potential. The d-term has two parts: one part from terms like $\left\langle d_{2}^{3}\right\rangle d_{2}^{3}$, and the other part where each $d_{2}^{A}$ in $\tfrac{1}{2}d_{2}^{A}d_{2}^{A}$ has one vev and one Higgs field. Those d-term contributions can be found by first re-expanding the d term part of the scalar potential of eq (2.1) using the following basis of diagonal U(3)$\times$×U(3) generators:
@@
$i\phi_{21}=\left(\bar{\phi}_{21}+{\tfrac{1}{\sqrt{2}}}h_{21}\right)\left(1,0,0\right)^{T}+...,$ (3.9)
where $h_{21}$ is the real scalar Higgs boson field and ++...+ … includes the fields that get “eaten” by the W and Z gauge bosons.
@@
where it has been assumed that $g_{+}=g$ at the unification scale and that $g_{Y^{\prime}}$ is negligible due to reasons discussed below. The approximation has also assumed that $\bar{\phi}^{2}_{21}\gg\bar{\phi}_{22},\bar{\varphi}_{2}$.
@@
Although the theory considered here is not invariant to superspace translations, it is invariant to superspace gauge transformations that change scalars into fermions. Also, the theory’s unification-scale symmetry breaking leaves the superpotential almost equal to zero, with variations from zero only showing up at the electroweak scale. Parallels with supersymmetry then open the door to the possibility that quadratic divergences naturally cancel in this model down to the electroweak scale. Further work would be required to establish this.
The $h_{21}$ real scalar field described above has the same interactions with the $W$ boson, $Z$ boson, and top quark as does the Standard Model Higgs boson. So this model is consistent with measurements of Higgs boson decays and interactions involving these particles, since those measurements are consistent with the Standard Model.
On the other hand, interactions of $h_{21}$ with other quarks or leptons are different in this model than in the Standard Model. As described later in this section, leptons and quarks (other than the top) acquire masses from different mechanisms, not from Yukawa interactions with the Higgs Boson. Those particles do have interactions with $h_{21}$ via the nonperturbative terms described in the Appendix. More work would need to be done to determine whether those interactions are consistent with observations.
@@ 3.3 Quark masses and mixing
where the gaugino coupling $\hat{g}$ is discussed below in eq (3.14). A tilde is put on tree-level masses and superpotential couplings to show that they include quantum modifications from the Appendix. For all of the fermion mass matrices in this paper (including the above matrix) the rows have an upper undotted spin index (using the convention of [6]) while the columns have a lower undotted spin index (e.g. $\tilde{u}^{\left(1\right)\alpha}u_{W\alpha}^{G}$). Those indices are summed over, and they are suppressed.
@@
$m_{t}\simeq\hat{g}\bar{\phi}_{21}\,\,\textrm{Top quark}.$ (3.13)
@@
The first- and second-generation quark interactions with the $W$ boson, however, differ from those of the Standard Model. For the second-generation up-type quark (mostly charm), both its left- and right-chiral components interact with the $W$ boson. For the first-generation up-type quark (mostly up), only its right-chiral component interacts with the $W$ boson. This chiral flipping of the first generation is not ruled out by experiment, since the spin of the proton (and other light hadrons) comes primarily from gluons and orbital angular momentum [19]. As a result, it is not possible to experimentally disentangle the spins of up and down quarks from other hadronic spins and thereby verify that only left-handed up and down quarks interact with the $W$ boson. Of course, a requirement of this model is that it must be able to reproduce CKM data. It does this even better than the Standard Model, as described below.
Above in eq (3.5), it was shown that the magnitude of the vev $\bar{\phi}_{21}$ is approximately determined by the $W$ boson mass and the SU(2) weak coupling constant $g_{2}$. In this model, the top quark mass is also determined by $\bar{\phi}_{21}$, but multiplied by the gaugino coupling $\hat{g}$ rather than the weak coupling $g_{2}$ (at the unification scale, these are the same). To be consistent with observation, this model would need to show that the difference in the coupling constants $g_{2}$ vs. $\hat{g}$ at the electroweak scale is the amount needed for the model to correctly reproduce both the $W$ boson mass and the top quark mass.
To get the actual value for this model’s top quark mass, the calculation should be performed to determine how $\hat{g}$ runs in this model as the scale is lowered from the unification scale. Such a calculation is outside the scope of this paper.
@@
${\rm Rows:}\,\tilde{d}^{(2)},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{A}\,\,\,{\rm Columns:}\,d_{W}^{G},d^{A},d^{(2)},d_{W}^{A}$ (3.15)
$M_{D}=\left(\begin{array}[]{cccc}{\hat{g}\bar{\tilde{\phi}}_{22}}&{...}&{\tilde{m}_{222}}&{{\tilde{\Gamma}_{222}\bar{\phi}_{22}}}\\ {...}&{\tfrac{1}{\sqrt{2}}\tilde{\Gamma}_{\Phi}c\bar{\varphi}_{2}}&{{\tilde{\Gamma}_{222}\bar{\tilde{\phi}}_{22}}}&{\tilde{m}^{\prime}_{\Phi}}\\ {\tilde{m}^{\prime}_{\lambda}}&{\hat{g}\bar{\varphi}_{2}}&{\hat{g}\bar{\phi}_{22}}&{...}\\ {\hat{g}c\bar{\varphi}_{2}}&{\tilde{m}^{\prime\prime}_{\Phi}}&{...}&{\tfrac{1}{\sqrt{2}}\tilde{\Gamma}_{\Phi}\bar{\varphi}_{2}}\end{array}\right),$
where only tree-level couplings are shown explicitly. Tildes again indicate that these couplings can be modified by quantum-generated couplings discussed in the appendix. Quantum couplings can also be generated for the components without tree-level contributions; those are denoted by “...”. Primes and double primes on $\tilde{m}^{\prime}_{\Phi}$ and $\tilde{m}^{\prime}_{\lambda}$ are used to show that these masses in the down-type matrix do not need to be the same as those in the up-type matrix, as discussed in the Appendix.
To a first approximation, the masses of the $d$, $s$, and $b$ quarks as well as an additional $f$ quark are given by the diagonal elements of eq (3.15). For example, $m_{b}\simeq\hat{g}\bar{\phi}_{22}$. Comparison with the top quark mass sets the value for $\bar{\phi}_{22}$ at $\bar{\phi}_{22}\simeq\bar{\phi}_{21}m_{b}/m_{t}$. The $f$ quark mass is primarily determined by the adjoint vev $\varphi_{2}$. In this model, none of the down-type quarks get their masses from the Higgs vev $\bar{\phi}_{21}$.
For this model, the numerical values of the quark mass matrices (in GeV) are assumed to be:
@@
$\mathit{M}_{U}=\left(\begin{array}[]{ccc}{-0.0033}&{0.0000}&{-0.0277}\\ {0.1532}&{1.2316}&{0.9688}\\ {7.2510}&{-7.8110}&{171.59}\end{array}\right).$
@@
$\mathit{M}_{D}=\left(\begin{array}[]{cccc}{0.0039}&{-0.0232}&{0.1192}&{0.0503}\\ {-0.0330}&{0.0872}&{0.1953}&{-0.1188}\\ {0.0093}&{-0.0589}&{4.1507}&{0.2934}\\ {0.2497}&{0.0127}&{-0.4219}&{2.8693}\end{array}\right).$
@@
The numerical values of these quark mass matrices were chosen by a fit to quark masses and CKM data, while not being in conflict with $Z$, $Z^{\prime}$, and exotic hadron data. In section 5, there is a discussion about how the theory is able to reproduce the many experiments that would seemingly rule out the possibility of an additional light quark. As more data is collected, it may be that rotations of these quark matrices that still reproduce quark masses and CKM data may provide a better global fit. Future work should also verify that there are no inconsistencies between these matrices and the quantum calculations described in the Appendix.
@@
$V_{L}^{U}\mathit{M}_{U}V_{R}^{U{\dagger}}={\rm diag}(-m_{u},m_{c},m_{t})$ (3.19)
$V_{L}^{D}\mathit{M}_{D}V_{R}^{D{\dagger}}={\rm diag}(-m_{d},m_{s},m_{b},m_{f})\,,$
@@
±$\pm V_{L}^{U}\left(\begin{array}[]{cccc}{0}&{0}&{0}&{0}\\ {0}&{1}&{0}&{0}\\ {0}&{0}&{1}&{0}\end{array}\right)V_{L}^{D{\dagger}}$
@@
$\left|V_{\rm{CKM}}^{+}\right|=\left(\begin{array}[]{cccc}{0.9737}&{0.2230}&{0.0039}&{0.0013}\\ {0.2241}&{0.9718}&{0.0394}&{0.9569}\\ {0.0072}&{0.0393}&{0.9934}&{0.0619}\end{array}\right)$ (3.23)
$\left|V_{\rm{CKM}}^{-}\right|=\left(\begin{array}[]{cccc}{0.9732}&{0.2253}&{0.0043}&{0.0012}\\ {0.2183}&{0.9750}&{0.0430}&{1.0414}\\ {0.0824}&{0.0188}&{0.9936}&{0.1451}\end{array}\right).$
@@
$\begin{array}[]{|c|c|c|}\hline\cr\bar{u}_{i}\gamma^{\mu}u_{i}Z_{\mu}&g_{L}&g_{R}\\ \hline\cr u^{\prime}&-\tfrac{2}{3}x&\tfrac{1}{2}-\tfrac{2}{3}x\\ c^{\prime}&\tfrac{1}{2}-\tfrac{2}{3}x&\tfrac{1}{2}-\tfrac{2}{3}x\\ t^{\prime}&\tfrac{1}{2}-\tfrac{2}{3}x&-\tfrac{2}{3}x\\ \hline\cr{\rm SM}&\tfrac{1}{2}-\tfrac{2}{3}x&-\tfrac{2}{3}x\\ \hline\cr\end{array}$ (3.24)
@@
$=\left(\begin{array}[]{cccc}{0.0520}&{-0.1082}&{0.0141}&{0.0066}\\ {-0.1082}&{-0.3970}&{-0.0056}&{0.0221}\\ {0.0141}&{-0.0056}&{-0.4172}&{-0.0505}\\ {0.0066}&{0.0221}&{-0.0505}&{0.0709}\end{array}\right),$

### v9

@@ 2 Dynamical Symmetry Breaking
$T_{23}^{Y^{\prime}}=T_{22}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(1,1,1,-1,-1,0\right)$ (2.15)
$T_{21}^{Y^{\prime}}=T_{1F}^{Y^{\prime}}$ $=\tfrac{1}{\sqrt{10}}{\rm diag}\left(-2,-2,-2,1,1,2\right)$
When acting on gaugino or adjoint fields, $T_{22}^{Y^{\prime}}$ should be used, and either version of $T_{mF}^{Y}$ can be used (since $T_{+}^{0}$ commutes with gauginos and adjoint fields).
It will be seen below that the $T_{mF}^{Y}$ matrices have the correct form for their gauge boson $A_{\mu}^{Y}$ to be identified as the U(1) weak hypercharge field of the Standard Model with ${\tfrac{1}{2}}\sqrt{{\tfrac{3}{5}}}g_{Y}$ identified as the weak hypercharge coupling. The $A_{\mu}^{Y^{\prime}}$ gauge field with its coupling $g_{Y^{\prime}}$ is a second U(1) gauge boson in this model that remains massless at the unification scale. Due to eq (2.10), there is no mixing between the $A_{\mu}^{Y}$ and $A_{\mu}^{Y^{\prime}}$ gauge fields. Below the unification scale, the couplings for the SU(3), SU(2), $Y$ and $Y^{\prime}$ groups run differently, so they are denoted by $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$.
@@
The angle $\theta_{Z}$ is chosen to make $(T_{21}^{\gamma})_{44}=(T_{22}^{\gamma})_{55}=0$, so that the photon gets no mass from $\bar{\phi}_{21}$, $\bar{\varphi}_{2}$, or any of the quantum-generated scalar vevs discussed in section 3.
The angle $\phi_{Z}$ is chosen to make $(T_{21}^{Z^{\prime}})_{44}=0$, so that it gets no mass from $\bar{\phi}_{21}$. The resulting angles are given by:
@@
$\tan\phi_{Z}$ $=-\tfrac{2}{\sqrt{10}}\cos\theta_{Z}g_{Y^{\prime}}/g_{2}\,.$
In section 4, it will be argued that $g_{Y^{\prime}}$ will be driven to a very small value from an effective anomaly. As a result, the angle $\phi_{Z}$ is very small, and $\theta_{Z}$ is very close to the Weinberg angle $\theta_{W}$ of the Standard Model.
@@
In the above labelling, lower numerical indices are fundamental-representation indices for the unbroken SU(3) group (the strong interaction). Fermions with a “WW” index interact with the $W$ boson (as members of an isodoublet). Based on the magnitude of their electric charges, up-type quarks, down-type quarks, charged leptons, and neutrinos are labelled with $u,d,e,\nu$. All of the fermion fields are 2-component Weyl fermions with a lower, undotted spin index. In the convention of [6] (which is also the Wess/Bagger & Bilal convention of [14]), a Weyl fermion with a lower undotted index corresponds to a right-chiral fermion that vanishes when acted on by $1-\gamma_{5}$ (see appendix A of [15]).
In that convention, the $u$ and $d$ fields are right-chiral fermions with electric charges of $\tfrac{2}{3}$} and $\tfrac{1}{3}$}, respectively, so they are mapped to right-chiral quarks. The fields $\tilde{u}$ and $\tilde{d}$ are right-chiral fermions with electric charges of $\tfrac{2}{3}$} and $\tfrac{1}{3}$}, respectively, so they are mapped to Hermitian conjugates of left-chiral quarks. One way that this model differs from the Standard Model is that some of the right-chiral quarks have a “WW” index so they interact with the $W$ boson, while some of the left-chiral quarks lack that index so they do not interact with the $W$ boson. That difference is discussed in the next section.
@@ 3.1 Confinement
$\sqrt{2}g_{3}i\int d^{4}x\rm{tr}\left(\tilde{\chi}\left\langle\varphi_{1}^{\dagger}\right\rangle\lambda-\tilde{\lambda}\left\langle\varphi_{1}^{\dagger}\right\rangle\chi\right)+h.c.,$ (3.2)
@@
There is another quantum effect pertinent to this section. For supersymmetric theories, instantons can generate nonperturbative low-energy effective superpotential terms. The appendix proposes that a similar effect occurs for this theory. To accommodate these terms, the following vevs that are zero classically acquire small but nonzero vevs quantum mechanically:
$\bar{\phi}_{22},\bar{\tilde{\phi}}_{21},\bar{\tilde{\phi}}_{22}\neq 0\textrm{ quantum mechanically}.$ (3.3)
A quantum superpotential term proportional to $1/\bar{\tilde{\phi}}_{21}^{2}$ (see eq (A.13)) sets the scale for masses of the heavy right-handed neutrinos discussed later in this section.
@@ 3.2 Observed boson masses
Phenomenologically, the mass of the $W$ boson determines the vacuum expectation value (vev) $\bar{\phi}_{21}$, with the running coupling $g_{2}^{2}$ evaluated at the $W$ boson mass scale.
The $Z$ boson mass in this model is also primarily determined by $\bar{\phi}_{21}$. Due to the form of eq (2.16), the $Z$ boson mass in this model differs slightly from the Standard Model expression. The mass of the $Z$ boson is:
@@
[…]
@@
$i\phi_{21}=\left(\bar{\phi}_{21}+{\tfrac{1}{\sqrt{2}}}h_{t}\right)\left(1,0,0\right)^{T}+$..., (3.8)
where $h_{t}$ is a real scalar, and ++... includes the fields that get “eaten” by the $W$ and $Z$ gauge bosons.
@@
where it has been assumed that $g_{+}=g$ at the unification scale and that $g_{Y^{\prime}}$ is negligible due to reasons discussed below. The approximation has also assumed that $\bar{\phi}^{2}_{21}\gg\bar{\varphi}_{2}$.
@@
Although the theory considered here is not invariant to superspace translations, it is invariant to superspace gauge transformations that change scalars into fermions. Also, the theory’s unification-scale symmetry breaking leaves the superpotential almost equal to zero, with variations from zero only showing up at the electroweak scale. Parallels with supersymmetry then open the door to the possibility that quadratic divergences naturally cancel in this model down to the electroweak scale. Further work would be required to prove this.
The $h_{t}$ real scalar field described above has the same interactions with the $W$ boson, $Z$ boson, and top quark as does the Standard Model Higgs boson. So this model is consistent with measurements of Higgs boson decays and interactions involving these particles, since those measurements are consistent with the Standard Model.
On the other hand, leptons and quarks (other than the top) have Higgs-like interactions with a different scalar $(\varphi_{2})_{23}$, acquiring their masses from its vev $\bar{\varphi}_{2}$. Nonetheless, as described in section 5, this model predicts results for almost all Higgs measurements that match those of the Standard Model, and therefore also match data. One area of difference is that this model predicts significantly smaller cross sections than the Standard Model for ttH production followed by $H\to b\bar{b}$ decay. Current data are not inconsistent with that prediction.
@@ 3.3 Quark masses and mixing
where the gaugino coupling $\hat{g}$ is discussed below in eq (3.13). A tilde is put on tree-level masses and superpotential couplings to show that they include quantum modifications such as those discussed in the appendix. For all of the fermion mass matrices in this paper (including the above matrix) the rows have an upper undotted spin index (using the convention of [6]) while the columns have a lower undotted spin index (e.g. $\tilde{u}^{\left(1\right)\alpha}u_{W\alpha}^{G}$). Those indices are summed over, and they are suppressed.
@@
$m_{t}\simeq\hat{g}\bar{\phi}_{21}\,\,\,\,\textrm{Top quark ass}.$ (3.12)
@@
The first- and second-generation quark interactions with the $W$ boson, however, differ from those of the Standard Model. For the second-generation up-type quark (mostly charm), both its left- and right-chiral components interact with the $W$ boson. For the first-generation up-type quark (mostly up), only its right-chiral component interacts with the $W$ boson. This chiral flipping of the first generation is not ruled out by experiment, since the spin of the proton (and other light hadrons) comes primarily from gluons and orbital angular momentum [19]. As a result, it is not possible to experimentally disentangle the spins of up and down quarks from other hadronic spins and thereby verify that only left-handed up and down quarks interact with the $W$ boson. This is discussed further in section 5.
In eq (3.4), it was shown that the $W$ boson mass is approximately determined by the vev $\bar{\phi}_{21}$ and the SU(2) weak coupling constant $g_{2}$. In this model, the top quark mass is also determined by $\bar{\phi}_{21}$, but multiplied by the gaugino coupling $\hat{g}$ rather than the weak coupling $g_{2}$ (at the unification scale, these are the same).
To get this model’s prediction for the top quark mass, the calculation should be performed to determine how $\hat{g}$ runs in this model as the scale is lowered from the unification scale. Such a calculation is outside the scope of this paper.
@@
${\rm Rows:}\,\tilde{d}^{(2)},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{A}\,\,\,{\rm Columns:}\,d_{W}^{G},d^{(2)},d^{A},d_{W}^{A}$ (3.14)
$M_{D}=\left(\begin{array}[]{cccc}{\hat{g}\bar{\tilde{\phi}}_{22}}&{\tilde{m}_{222}}&{...}&{{\tilde{\Gamma}_{222}\bar{\phi}_{22}}}\\ {...}&{{\tilde{\Gamma}_{222}\bar{\tilde{\phi}}_{22}}}&...&{\tilde{m}^{\prime}_{\Phi}}\\ {\tilde{m}^{\prime}_{\lambda}}&{\hat{g}\bar{\phi}_{22}}&{\hat{g}\bar{\varphi}_{2}}&{...}\\ {...}&{...}&{\tilde{m}^{\prime\prime}_{\Phi}}&{\tfrac{1}{\sqrt{2}}|\tilde{\Gamma}_{\Phi}|\bar{\varphi}_{2}}\end{array}\right),$
where only tree-level couplings are shown explicitly. Tildes again indicate that these couplings can be modified by quantum-generated couplings discussed in the appendix. Quantum couplings can also be generated for the components without tree-level contributions; those are denoted by “...”. Primes and double primes on $\tilde{m}^{\prime}_{\Phi}$ and $\tilde{m}^{\prime}_{\lambda}$ are used to show that these masses in the down-type matrix do not need to be the same as those in the up-type matrix.
To a first approximation, the masses of the $d$, $s$, and $b$ quarks as well as an additional $f$ quark are given by the diagonal elements of eq (3.14). For example, $m_{b}\simeq\hat{g}\bar{\varphi}_{2}$. Comparison with the top quark mass sets the value for this vev at $\bar{\varphi}_{2}\simeq(m_{b}/m_{t})\bar{\phi}_{21}$. The $f$ quark mass is also generated by $\bar{\varphi}_{2}$, but with a superpotential coupling. None of the down-type quarks get their masses from the vev $\bar{\phi}_{21}$ that gives mass to the top quark.
One model for the numerical values of the quark mass matrices (in GeV) is to take them to be the following:
@@
$\mathit{M}_{U}=\left(\begin{array}[]{ccc}{0}&{0.0198}&{0.7302}\\ {-0.1309}&{1.2736}&{0}\\ {-0.4147}&{4.0973}&{171.95}\end{array}\right).$
@@
$\mathit{M}_{D}=\left(\begin{array}[]{cccc}{.0021}&{-0.0164}&{-0.0664}&{-0.0193}\\ {0.0303}&{0.0759}&{0.1792}&{-0.0788}\\ {-0.0414}&{-0.3295}&{4.1217}&{0.3959}\\ {0.2953}&{-0.0532}&{0.5825}&{-2.8053}\end{array}\right).$
@@
The numerical values of these quark mass matrices were chosen by a fit to quark masses and CKM data, while not being in conflict with $Z$, $Z^{\prime}$, and exotic hadron data. The two elements in $\mathit{M}_{U}$ that are proportional to $\bar{\tilde{\phi}}_{21}$ were set to zero to make them compatible with the large neutrino masses discussed in eq (A.13) of the appendix. In section 5, there is a discussion about how the theory is able to reproduce the many experiments that would seemingly rule out the possibility of an additional light quark. As more data is collected, it may be that rotations of these quark matrices that still reproduce quark masses and CKM data may provide a better global fit.
@@
$V_{L}^{U}\mathit{M}_{U}V_{R}^{U{\dagger}}={\rm diag}(m_{u},m_{c},m_{t})$ (3.18)
$V_{L}^{D}\mathit{M}_{D}V_{R}^{D{\dagger}}={\rm diag}(m_{d},m_{s},m_{b},m_{f})\,$,
@@
$\pm V_{L}^{U}\left(\begin{array}[]{cccc}{0}&{0}&{0}&{0}\\ {0}&{1}&{0}&{0}\\ {0}&{0}&{1}&{0}\end{array}\right)V_{L}^{D{\dagger}}$
@@
$\left|V_{\rm{CKM}}^{+}\right|=\left(\begin{array}[]{cccc}{0.9740}&{0.2231}&{0.0040}&{0.0021}\\ {0.2240}&{0.9732}&{0.0394}&{1.0282}\\ {0.0079}&{0.0394}&{0.9890}&{0.1183}\end{array}\right)$ (3.22)
$\left|V_{\rm{CKM}}^{-}\right|=\left(\begin{array}[]{cccc}{0.9749}&{0.2261}&{0.0046}&{0.0032}\\ {0.2182}&{0.9747}&{0.0429}&{0.9712}\\ {0.0077}&{0.0394}&{0.9891}&{0.1662}\end{array}\right).$
@@
[…]
@@
$=\left(\begin{array}[]{cccc}{-0.0941}&{-0.0743}&{-0.0058}&{-0.0018}\\ {-0.0743}&{-0.4056}&{-0.0004}&{-0.0115}\\ {-0.0058}&{-0.0004}&{-0.4159}&{0.0482}\\ {-0.0018}&{-0.0115}&{0.0482}&{-0.0844}\end{array}\right),$